OpenStudy (anonymous):

Can you help me solve this? I'm not quite sure how to do it. I have two equations: n+q=9 (0.05)n+(0.25)q=1.05

OpenStudy (austinl):

Ok, we can either do substitution which may be more involved. Or we can do elimination. What would you care for? \[n+q=9\] \[0.05n+0.25q=1.05\]

OpenStudy (anonymous):

either is fine(:

OpenStudy (austinl):

Let's do elimination. First we we should choose a variable to eliminate first. Let's go with x. We would multiply the top equation by -0.05 \[-0.05\times(n+q=9)\]= \[-0.05n-0.05q=-0.45\] and we still have the second equation \[0.05n+0.25q=1.05\]

OpenStudy (austinl):

We would add them together top to bottom. Can you do that for me?

OpenStudy (anonymous):

Okay. I'm a bit confused on how to add them though.

OpenStudy (anonymous):

If I set it up I have this:  but what do I do where there is a n+q?

OpenStudy (austinl):

 Then you simply solve for q!

OpenStudy (austinl):

Then we can plug that value into one of the other equations.

OpenStudy (anonymous):

Oh, okay.

OpenStudy (austinl):

Did what I just did make sense to you?

OpenStudy (anonymous):

yes, but I assume that the 0.20q+0.60 is supposed to be 0.20q=0.60?

OpenStudy (austinl):

Yeah, sorry! Silly typo.

OpenStudy (anonymous):

Okay! And then we just fill in q in the other equation and solve?

OpenStudy (austinl):

Correct, what did you get as a value for q?

OpenStudy (anonymous):

3

OpenStudy (austinl):

Correct! So when you plug it into the first equation we were given. \[n+q=9\] \[n+3=9\] n=?

OpenStudy (anonymous):

6

OpenStudy (austinl):

Correct! n=6 q=3

OpenStudy (anonymous):

Great, thank you so much! I've been having a bit of a hard time understanding how to do these types of problems, and you cleared it up!

OpenStudy (austinl):

No problem! I love algebra, it is fun to me however geeky that may sound lol If you have any other questions feel free to tag me in a question and I will do my best to help!

OpenStudy (anonymous):

Great, thank you so much! (: